# BevSplat: Resolving Height Ambiguity via Feature-Based Gaussian Primitives for Weakly-Supervised Cross-View Localization

### v3

@@ 4.1 Comparison with State-of-the-Art Methods
We compare our method with the latest state-of-the-art (SOTA) approaches, including supervised methods such as Boosting [4], VFA [20], CCVPE [57], HC-Net [6], and DenseFlow [18], all of which rely on ground-truth camera poses for supervision. We also compare with G2Sweakly [1], which uses only a satellite image and a corresponding ground image as input, similar to our setup.
KITTI. The comparison results on the KITTI dataset [51] are summarized in Table 1. Since our rotation estimator is inherited from G2Sweakly [1], the rotation estimation performance is identical between the two methods. However, our method significantly outperforms G2Sweakly [1] in terms of location estimation across almost all evaluation metrics, yielding substantial improvements in both longitudinal pose accuracy and the corresponding mean and median errors. This improvement can be attributed to the limitations of the IPM projection method used in G2Sweakly [1], which suffers from distortions in scenes that are far from the camera and fails to capture the details of objects above the ground plane.
Our feature-based Gaussian splatting for BEV synthesis effectively addresses these issues, leading to a notable enhancement in localization accuracy. Fig. 1 and Fig. 4 visualize the difference between the IPM projection and our proposed BEV synthesis method, clearly demonstrating that our projection technique resolves challenges such as occlusions caused by tall objects (e.g., buildings, trees, vehicles) and geometric distortions from curved roads. Furthermore, in cross-area evaluations, our method even surpasses supervised approaches (Boosting [4], VFA [20], CCVPE [57], HC-Net [6], and DenseFlow [18]) in terms of mean and median errors, showcasing the strong generalization ability of our approach and highlighting the potential of weakly supervised methods.
It is worth noting that the experimental results for VFA [20] are taken from the PIDLoc [22] paper. This is because the authors of VFA have adopted a setting in their prior works [58, 59] that aligns ground-truth poses to the satellite image center, which risks overfitting by biasing predictions towards the center. Since the official code for VFA [20] is not available, we attempted to reproduce their results and found that we could only achieve the reported performance by using this same setting. However, when we use the standard setting adopted by our method and other comparable works [4, 57, 6, 18, 1], our reproduced results are consistent with the VFA results reported in PIDLoc [22]. Therefore, for a fair comparison, we use the VFA [20] results from PIDLoc [22].
@@
Table 2: Comparison with the most recent state-of-the-art (SOTA) on VIGOR (* denotes supervised learning algorithms). Our weakly supervised approach slightly outperforms SOTA supervised methods in Same-Area evaluations and comprehensively surpasses them in Cross-Area evaluations.
@@
VIGOR. The comparison results on the VIGOR dataset [15] are presented in Table 2. Our method demonstrates a significant reduction in both mean and median errors compared to the baseline weakly supervised approach, G2SWeakly [1], across all evaluation scenarios. Furthermore, even when benchmarked against the state-of-the-art method in fully supervised settings, our method achieves notable improvements on the majority of metrics in both same-area and cross-area evaluations. This reduces the gap between weakly supervised and fully supervised methods, indicating that our approach generalizes effectively to diverse localization tasks, including both same-area and cross-area scenarios, as well as cases where the query images are either panoramic or captured using pinhole cameras.
Computation comparison. All evaluations of GPU memory usage were performed on an NVIDIA RTX 4090. Our model, which features a DINOv2 [49] backbone fine-tuned with our lightweight DPT network [50] (composed of a few CNN layers), requires considerably less memory during the training phase compared to G2SWeakly [1]. On the KITTI [51] dataset (batch size 8), our training phase uses only 9.2 GB of GPU memory, substantially less than the 22.7 GB required by G2SWeakly [1]. During inference, our model consumes 7.7 GB of GPU memory, while G2SWeakly [1] requires 7.2 GB as shown in Table 5. This increase 0.5 GB for our method is attributable to the larger DINOv2 backbone [49], representing a trade-off for its enhanced feature representation capabilities.
@@
Lift-Splat-Shoot [60] 1 16.14 13.94 17.74 14.51
OrienterNet [5] 0 15.59 13.68 16.15 13.80
@@
Lift-Splat-Shoot [60] 1 7.89 4.30 11.63 5.31
OrienterNet [5] 1 5.71 3.20 10.02 5.07
@@
OrienterNet [5] $32.4$ $10.8$ $71$
LSS [60] $26.1$ $8.3$ $85$
G2SWeakly [1] $22.7$ $7.2$ $31$
@@ 4.2 Ablation Study
Different BEV synthesis approaches. To validate the effectiveness of our BevSplat method, we compared it against two common BEV generation techniques: the Inverse Perspective Mapping (IPM) approach as utilized in [1], and direct projection of 3D point clouds. For a fair comparison, both baseline methods and our BevSplat employed the same DINOv2 backbone for feature extraction.
@@
Direct point cloud projection to BEV. While this approach can address some of the aforementioned IPM limitations, it introduces new challenges stemming from point cloud sparsity (visualized in Fig. 4(c)). This sparsity leads to BEV voids and discontinuous features, exacerbated by the lack of control over individual point attributes such as opacity, scale, and shape. Furthermore, unlike 3DGS [7] which utilizes $\alpha$-blending, the simple top-down projection inherent in this method causes severe occlusion, leading to the loss of underlying feature information—an issue also evident in Fig. 4(c). Our ablation study (Table 5) confirms this inherent characteristic: BevSplat(w/o OPT) configured with non-optimizable Gaussian parameters (e.g., fixed opacity=1.0, scale=0.1, offsets=0, and no learned adjustments) performs comparably to direct point cloud projection, underscoring that point clouds can be seen as a degenerate form of 3DGS [7]. However, our full BevSplat(w/ OPT) formulation significantly improves upon this baseline by optimizing Gaussian opacity, scale, shape, and position. Guided by satellite imagery, this optimization process effectively mitigates the issues of point cloud projection, yielding coherent, feature-rich BEV representations and thereby enabling superior localization accuracy. As detailed in Table 5, BevSplat subsequently outperforms both IPM and Direct Projection methods.
Other depth-based re-sampling methods. We further compare our method with other BEV projection techniques that are also based on depth prediction, such as those used in Lift-Splat-Shoot [60] and OrienterNet [5] as detailed in Table 5, by adapting their projection modules into our framework. Although these approaches can generate a denser BEV and, like our method, leverage the full vertical information from the ground-view image—allowing them to perform well in same-area settings when guided by GPS labels—they fail to generalize to unseen environments. They tend to make erroneous guesses to fill in occluded regions, which explains their performance degradation in cross-area evaluations. Furthermore, they employ a simple weighted averaging for BEV projection, which is less accurate for handling vertical occlusions compared to BevSplat’s principled alpha blending. Finally, both methods utilize a complex $h\times w\times d$ depth representation to perform an attention-based sum over ground features. This implicit, high-dimensional process incurs substantial computational and memory overhead to produce a denser BEV as shown in Table 5.
@@
Impact of Different Foundation Models and Weights. Prioritizing robust outdoor generalization and effective 3D-relevant feature extraction for our foundation model, we selected DINOv2 [49] fine-tuned with the FiT method [48]. This model, which we term DINOv2(FiT), utilizes its dinov2_base_fine pre-trained weights renowned for these capabilities. To validate this choice and compare its efficacy against alternatives, our ablation study also evaluated VGG [62], DINOv1 [63], and the original DINOv2. All DINO-based backbones in this study (DINOv1, DINOv2, and DINOv2(FiT)) were subsequently further fine-tuned by us using a DPT-like module [50]. The ablation results (Table 6) validated our selection, as DINOv2(FiT), after our DPT-like fine-tuning, demonstrated superior performance among the evaluated backbones.
Impact of Different Fine-tuning Methods. Using DPT-like models is a common practice for 3D vision tasks; for example, VGGT [64] utilizes DPT [50] for point cloud reconstruction, depth estimation, and feature matching. Although we are the first to apply DPT-DINO for feature extraction in the specific sub-field of cross-view localization, our motivation is to similarly obtain features that are rich in 3D information. This is analogous to human navigation, where in addition to semantic information, an understanding of the real 3D scene is also crucial for localization. However, obtaining such 3D-aware features to bridge the significant ground-satellite domain gap is non-trivial. Simpler fine-tuning methods like direct end-to-end training or LoRA [61] fail, as they either lose crucial texture details or are not powerful enough to adapt the foundation model. We use a DPT-like module because its multi-scale feature fusion architecture is uniquely suited for this challenge. It successfully adapts the backbone by preserving both the low-level texture and high-level semantic information required for matching across these different domains as shown in Table 5.
@@ 4.3 Multi-Frame Localization
Beyond processing single ground-level images, our method extends to leveraging multiple frames from video sequences to enhance localization robustness, particularly in dynamic environments. Similar to CVLNet [65], given known inter-frame relative poses, our BevSplat technique projects features from these frames into a unified BEV. These multi-frame BEV features are then fused by a Transformer employing self-attention. We investigated this multi-frame capability using query video sequences comprising 1 to 6 frames. The results, presented in Fig. 6, demonstrate that localization performance consistently improves with an increasing number of frames in the sequence. This underscores the efficacy of our approach in leveraging temporal information from video data for enhanced localization robustness.
@@ 5 Conclusion
The authors are grateful for the valuable comments and suggestions by the reviewers and ACs. This work was supported by NSFC (62406194), Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI), MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (KLIP-HuMaCo). A part of the experiments of this work were supported by the core facility Platform of Computer Science and Communication, SIST, ShanghaiTech University.
@@ Appendix A Carefule Explanation of the Weakly Supervised Setup
We focus on a weakly-supervised setting because precise GPS data is often unavailable in the real world. To do this, we adopt the weakly-supervised setup from G2SWeakly [1], which defines two scenarios:
@@ Appendix D Robustness to Localization Errors
Location
Error (m2)
$\boldsymbol{\lambda_{1}}$ Same Area Cross Area
@@ Appendix F Ablation Study on the Number of Gaussian Primitives Per Pixel($N_{p}$)
This finding is also consistent with prior work. Our design was inspired by PixelSplat [66], which similarly found Np=3 to be a robust and effective setting across multiple datasets. Therefore, we conclude that Np=3 is a well-justified hyperparameter that should be generally applicable.
@@ Appendix G Applicability to Multi-Frame Localization Tasks
As detailed in the main paper, our quantitative analysis of a CVLNet-based multi-frame fusion method [65] demonstrated progressively improved performance with an increasing number of frames, confirming its efficacy for temporal sequence tasks. To complement these findings, Figure 7 provides a qualitative illustration.
@@
Mean(m)
Median(m)
Mean(m)
Median(m)
G2SWeakly 6 0
8.65($\downarrow$ 4.1%)
5.22($\downarrow$ 5.7%)
9.41($\downarrow$ 4.6%)
6.01($\downarrow$ 5.3%)
Direct Projection 6 0
7.05($\downarrow$ 7.1%)
3.86($\downarrow$ 9.2%)
8.15($\downarrow$ 8.7%)
5.31($\downarrow$ 8.6%)
Ours 6 0
5.01($\downarrow$ 13.9%)
2.27($\downarrow$ 20.4%)
6.09($\downarrow$ 13.6%)
2.71($\downarrow$ 15.8%)
G2SWeakly 6 1
6.25($\downarrow$ 6.4%)
3.38($\downarrow$ 8.9%)
8.18($\downarrow$ 4.9%)
4.32($\downarrow$ 10.7%)
Direct Projection 6 1
3.85($\downarrow$ 13.1%)
2.91($\downarrow$ 11.6%)
7.18($\downarrow$ 9.6%)
3.96($\downarrow$ 14.7%)
Ours 6 1
2.01($\downarrow$ 30.0%)
1.77($\downarrow$ 14.1%)
5.23($\downarrow$ 15.6%)
1.94($\downarrow$ 22.7%)
@@ Appendix I Sensitivity to Depth Prediction Quality and Failure Cases
We evaluate our framework’s performance with three different depth foundation models: DepthAnythingv1 [67], ZoeDepth [68], and DepthAnythingv2 [69] in Table 11:
@@
DepthAnythingV1 [67] 0 $5.91$ $2.84$ $7.21$ $3.25$
ZoeDepth [68] 0 $5.84$ $2.86$ $7.14$ $3.22$
DepthAnythingV2 [69] 0 $5.82$ $2.85$ $7.05$ $3.22$
DepthAnythingV1 [67] 1 $2.97$ $2.11$ $6.28$ $2.52$
ZoeDepth [68] 1 $2.91$ $2.03$ $6.21$ $2.54$
DepthAnythingV2 [69] 1 $2.87$ $2.06$ $6.20$ $2.51$
@@
To test our method’s robustness, we generated synthetic Rain, Fog, and Night data for KITTI (following the methodology of Robust-Depth [70]). This allows for a controlled comparison against the G2SWeakly baseline under challenging conditions.
@@
Mean(m) $\downarrow$
Median(m) $\downarrow$
Mean(m) $\downarrow$
Median(m) $\downarrow$
@@
Our methodology employs a world coordinate system consistent with the OpenCV convention [71], as depicted in Figure 8. This is a right-handed system where, from the camera’s viewpoint, the +X axis extends to its right, the +Y axis points downwards, and the +Z axis aligns with its forward viewing direction. Consequently, the upward direction corresponds to the -Y axis.
@@
As discussed in the main paper’s conclusion, a current limitation of our BevSplat method, which renders Bird’s-Eye View (BEV) perspectives based on 3D Gaussian Splatting [7], is its computational speed compared to Inverse Perspective Mapping (IPM). For instance, on an NVIDIA RTX 4090 GPU, BevSplat requires 14 ms to generate a single BEV image. In contrast, IPM, which utilizes direct linear interpolation, can achieve this in 4ms. This performance disparity affects the overall inference speed of our model. Therefore, a significant direction for our future work is the exploration of faster and more compact Gaussian representations to address this bottleneck and enhance real-time applicability.

### v4

@@ 4.1 Comparison with State-of-the-Art Methods
We compare our method with the latest state-of-the-art (SOTA) approaches, including supervised methods such as Boosting [shi2023boosting], VFA [wang2024view], CCVPE [xia2023convolutional], HC-Net [wang2024fine], and DenseFlow [song2024learning], all of which rely on ground-truth camera poses for supervision. We also compare with G2Sweakly [shi2024weakly], which uses only a satellite image and a corresponding ground image as input, similar to our setup.
KITTI. The comparison results on the KITTI dataset [geiger2013vision] are summarized in Table 1. Since our rotation estimator is inherited from G2Sweakly [shi2024weakly], the rotation estimation performance is identical between the two methods. However, our method significantly outperforms G2Sweakly [shi2024weakly] in terms of location estimation across almost all evaluation metrics, yielding substantial improvements in both longitudinal pose accuracy and the corresponding mean and median errors. This improvement can be attributed to the limitations of the IPM projection method used in G2Sweakly [shi2024weakly], which suffers from distortions in scenes that are far from the camera and fails to capture the details of objects above the ground plane.
Our feature-based Gaussian splatting for BEV synthesis effectively addresses these issues, leading to a notable enhancement in localization accuracy. Fig. 1 and Fig. 4 visualize the difference between the IPM projection and our proposed BEV synthesis method, clearly demonstrating that our projection technique resolves challenges such as occlusions caused by tall objects (e.g., buildings, trees, vehicles) and geometric distortions from curved roads. Furthermore, in cross-area evaluations, our method even surpasses supervised approaches (Boosting [shi2023boosting], VFA [wang2024view], CCVPE [xia2023convolutional], HC-Net [wang2024fine], and DenseFlow [song2024learning]) in terms of mean and median errors, showcasing the strong generalization ability of our approach and highlighting the potential of weakly supervised methods.
It is worth noting that the experimental results for VFA [wang2024view] are taken from the PIDLoc [lee2025pidloc] paper. This is because the authors of VFA have adopted a setting in their prior works [wang2023view, wang2022satellite] that aligns ground-truth poses to the satellite image center, which risks overfitting by biasing predictions towards the center. Since the official code for VFA [wang2024view] is not available, we attempted to reproduce their results and found that we could only achieve the reported performance by using this same setting. However, when we use the standard setting adopted by our method and other comparable works [shi2023boosting, xia2023convolutional, wang2024fine, song2024learning, shi2024weakly], our reproduced results are consistent with the VFA results reported in PIDLoc [lee2025pidloc]. Therefore, for a fair comparison, we use the VFA [wang2024view] results from PIDLoc [lee2025pidloc].
@@
Table 2: Comparison with the most recent state-of-the-art (SOTA) on VIGOR (* denotes supervised learning algorithms). Our weakly supervised approach achieves performance comparable to SOTA supervised methods in Same-Area evaluations and comprehensively surpasses them in Cross-Area evaluations.
@@
VIGOR. The comparison results on the VIGOR dataset [zhu2021vigor] are presented in Table 2. Our method demonstrates a significant reduction in both mean and median errors compared to the baseline weakly supervised approach, G2SWeakly [shi2024weakly], across all evaluation scenarios. Furthermore, even when benchmarked against state-of-the-art fully supervised methods, our method maintains comparable performance in the same-area evaluations while achieving notable improvements across most metrics in cross-area evaluations. This reduces the gap between weakly supervised and fully supervised methods, indicating that our approach generalizes effectively to diverse localization tasks, including both same-area and cross-area scenarios, as well as cases where the query images are either panoramic or captured using pinhole cameras.
Computation comparison. All evaluations of GPU memory usage were performed on an NVIDIA RTX 4090. Our model, which features a DINOv2 [oquab2023dinov2] backbone fine-tuned with our lightweight DPT network [ranftl2021vision] (composed of a few CNN layers), requires considerably less memory during the training phase compared to G2SWeakly [shi2024weakly]. On the KITTI [geiger2013vision] dataset (batch size 8), our training phase uses only 9.2 GB of GPU memory, substantially less than the 22.7 GB required by G2SWeakly [shi2024weakly]. During inference, our model consumes 7.7 GB of GPU memory, while G2SWeakly [shi2024weakly] requires 7.2 GB as shown in Table 5. This increase 0.5 GB for our method is attributable to the larger DINOv2 backbone [oquab2023dinov2], representing a trade-off for its enhanced feature representation capabilities.
@@
Lift-Splat-Shoot [philion2020liftsplatshootencoding] 1 16.14 13.94 17.74 14.51
OrienterNet [sarlin2023orienternet] 0 15.59 13.68 16.15 13.80
@@
Lift-Splat-Shoot [philion2020liftsplatshootencoding] 1 7.89 4.30 11.63 5.31
OrienterNet [sarlin2023orienternet] 1 5.71 3.20 10.02 5.07
@@
OrienterNet [sarlin2023orienternet] $32.4$ $10.8$ $71$
LSS [philion2020liftsplatshootencoding] $26.1$ $8.3$ $85$
G2SWeakly [shi2024weakly] $22.7$ $7.2$ $31$
@@ 4.2 Ablation Study
Different BEV synthesis approaches. To validate the effectiveness of our BevSplat method, we compared it against two common BEV generation techniques: the Inverse Perspective Mapping (IPM) approach as utilized in [shi2024weakly], and direct projection of 3D point clouds. For a fair comparison, both baseline methods and our BevSplat employed the same DINOv2 backbone for feature extraction.
@@
Direct point cloud projection to BEV. While this approach can address some of the aforementioned IPM limitations, it introduces new challenges stemming from point cloud sparsity (visualized in Fig. 4(c)). This sparsity leads to BEV voids and discontinuous features, exacerbated by the lack of control over individual point attributes such as opacity, scale, and shape. Furthermore, unlike 3DGS [kerbl20233d] which utilizes $\alpha$-blending, the simple top-down projection inherent in this method causes severe occlusion, leading to the loss of underlying feature information—an issue also evident in Fig. 4(c). Our ablation study (Table 5) confirms this inherent characteristic: BevSplat(w/o OPT) configured with non-optimizable Gaussian parameters (e.g., fixed opacity=1.0, scale=0.1, offsets=0, and no learned adjustments) performs comparably to direct point cloud projection, underscoring that point clouds can be seen as a degenerate form of 3DGS [kerbl20233d]. However, our full BevSplat(w/ OPT) formulation significantly improves upon this baseline by optimizing Gaussian opacity, scale, shape, and position. Guided by satellite imagery, this optimization process effectively mitigates the issues of point cloud projection, yielding coherent, feature-rich BEV representations and thereby enabling superior localization accuracy. As detailed in Table 5, BevSplat subsequently outperforms both IPM and Direct Projection methods.
Other depth-based re-sampling methods. We further compare our method with other BEV projection techniques that are also based on depth prediction, such as those used in Lift-Splat-Shoot [philion2020liftsplatshootencoding] and OrienterNet [sarlin2023orienternet] as detailed in Table 5, by adapting their projection modules into our framework. Although these approaches can generate a denser BEV and, like our method, leverage the full vertical information from the ground-view image—allowing them to perform well in same-area settings when guided by GPS labels—they fail to generalize to unseen environments. They tend to make erroneous guesses to fill in occluded regions, which explains their performance degradation in cross-area evaluations. Furthermore, they employ a simple weighted averaging for BEV projection, which is less accurate for handling vertical occlusions compared to BevSplat’s principled alpha blending. Finally, both methods utilize a complex $h\times w\times d$ depth representation to perform an attention-based sum over ground features. This implicit, high-dimensional process incurs substantial computational and memory overhead to produce a denser BEV as shown in Table 5.
@@
Impact of Different Foundation Models and Weights. Prioritizing robust outdoor generalization and effective 3D-relevant feature extraction for our foundation model, we selected DINOv2 [oquab2023dinov2] fine-tuned with the FiT method [yue2025improving]. This model, which we term DINOv2(FiT), utilizes its dinov2_base_fine pre-trained weights renowned for these capabilities. To validate this choice and compare its efficacy against alternatives, our ablation study also evaluated VGG [simonyan2014very], DINOv1 [caron2021emerging], and the original DINOv2. All DINO-based backbones in this study (DINOv1, DINOv2, and DINOv2(FiT)) were subsequently further fine-tuned by us using a DPT-like module [ranftl2021vision]. The ablation results (Table 6) validated our selection, as DINOv2(FiT), after our DPT-like fine-tuning, demonstrated superior performance among the evaluated backbones.
Impact of Different Fine-tuning Methods. Using DPT-like models is a common practice for 3D vision tasks; for example, VGGT [wang2025vggtvisualgeometrygrounded] utilizes DPT [ranftl2021vision] for point cloud reconstruction, depth estimation, and feature matching. Although we are the first to apply DPT-DINO for feature extraction in the specific sub-field of cross-view localization, our motivation is to similarly obtain features that are rich in 3D information. This is analogous to human navigation, where in addition to semantic information, an understanding of the real 3D scene is also crucial for localization. However, obtaining such 3D-aware features to bridge the significant ground-satellite domain gap is non-trivial. Simpler fine-tuning methods like direct end-to-end training or LoRA [hu2021loralowrankadaptationlarge] fail, as they either lose crucial texture details or are not powerful enough to adapt the foundation model. We use a DPT-like module because its multi-scale feature fusion architecture is uniquely suited for this challenge. It successfully adapts the backbone by preserving both the low-level texture and high-level semantic information required for matching across these different domains as shown in Table 5.
@@ 4.3 Multi-Frame Localization
Beyond processing single ground-level images, our method extends to leveraging multiple frames from video sequences to enhance localization robustness, particularly in dynamic environments. Similar to CVLNet [shi2022cvlnet], given known inter-frame relative poses, our BevSplat technique projects features from these frames into a unified BEV. These multi-frame BEV features are then fused by a Transformer employing self-attention. We investigated this multi-frame capability using query video sequences comprising 1 to 6 frames. The results, presented in Fig. 6, demonstrate that localization performance consistently improves with an increasing number of frames in the sequence. This underscores the efficacy of our approach in leveraging temporal information from video data for enhanced localization robustness.
@@ 5 Conclusion
The authors are grateful for the valuable comments and suggestions by the reviewers and ACs. This work was supported by NSFC (62406194), Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI), MoE Key Laboratory of Intelligent Perception, HPC Platform of ShanghaiTech University and Human-Machine Collaboration (KLIP-HuMaCo). A part of the experiments of this work were supported by the core facility Platform of Computer Science and Communication, SIST, ShanghaiTech University.
@@ Appendix A Carefule Explanation of the Weakly Supervised Setup
We focus on a weakly-supervised setting because precise GPS data is often unavailable in the real world. To do this, we adopt the weakly-supervised setup from G2SWeakly [shi2024weakly], which defines two scenarios:
@@ Appendix D Robustness to Localization Errors
Location Error (m2) $\boldsymbol{\lambda_{1}}$ Same Area Cross Area
@@ Appendix F Ablation Study on the Number of Gaussian Primitives Per Pixel($N_{p}$)
This finding is also consistent with prior work. Our design was inspired by PixelSplat [charatan2024pixelsplat], which similarly found Np=3 to be a robust and effective setting across multiple datasets. Therefore, we conclude that Np=3 is a well-justified hyperparameter that should be generally applicable.
@@ Appendix G Applicability to Multi-Frame Localization Tasks
As detailed in the main paper, our quantitative analysis of a CVLNet-based multi-frame fusion method [shi2022cvlnet] demonstrated progressively improved performance with an increasing number of frames, confirming its efficacy for temporal sequence tasks. To complement these findings, Figure 7 provides a qualitative illustration.
@@
Mean(m) Median(m) Mean(m) Median(m)
G2SWeakly 6 0 8.65($\downarrow$ 4.1%) 5.22($\downarrow$ 5.7%) 9.41($\downarrow$ 4.6%) 6.01($\downarrow$ 5.3%)
Direct Projection 6 0 7.05($\downarrow$ 7.1%) 3.86($\downarrow$ 9.2%) 8.15($\downarrow$ 8.7%) 5.31($\downarrow$ 8.6%)
Ours 6 0 5.01($\downarrow$ 13.9%) 2.27($\downarrow$ 20.4%) 6.09($\downarrow$ 13.6%) 2.71($\downarrow$ 15.8%)
G2SWeakly 6 1 6.25($\downarrow$ 6.4%) 3.38($\downarrow$ 8.9%) 8.18($\downarrow$ 4.9%) 4.32($\downarrow$ 10.7%)
Direct Projection 6 1 3.85($\downarrow$ 13.1%) 2.91($\downarrow$ 11.6%) 7.18($\downarrow$ 9.6%) 3.96($\downarrow$ 14.7%)
Ours 6 1 2.01($\downarrow$ 30.0%) 1.77($\downarrow$ 14.1%) 5.23($\downarrow$ 15.6%) 1.94($\downarrow$ 22.7%)
@@ Appendix I Sensitivity to Depth Prediction Quality and Failure Cases
We evaluate our framework’s performance with three different depth foundation models: DepthAnythingv1 [depth_anything_v1], ZoeDepth [bhat2023zoedepthzeroshottransfercombining], and DepthAnythingv2 [yang2024depthv2] in Table 11:
@@
DepthAnythingV1 [depth_anything_v1] 0 $5.91$ $2.84$ $7.21$ $3.25$
ZoeDepth [bhat2023zoedepthzeroshottransfercombining] 0 $5.84$ $2.86$ $7.14$ $3.22$
DepthAnythingV2 [yang2024depthv2] 0 $5.82$ $2.85$ $7.05$ $3.22$
DepthAnythingV1 [depth_anything_v1] 1 $2.97$ $2.11$ $6.28$ $2.52$
ZoeDepth [bhat2023zoedepthzeroshottransfercombining] 1 $2.91$ $2.03$ $6.21$ $2.54$
DepthAnythingV2 [yang2024depthv2] 1 $2.87$ $2.06$ $6.20$ $2.51$
@@
To test our method’s robustness, we generated synthetic Rain, Fog, and Night data for KITTI (following the methodology of Robust-Depth [Gasperini_2023]). This allows for a controlled comparison against the G2SWeakly baseline under challenging conditions.
@@
Our methodology employs a world coordinate system consistent with the OpenCV convention [bradski2008learning], as depicted in Figure 8. This is a right-handed system where, from the camera’s viewpoint, the +X axis extends to its right, the +Y axis points downwards, and the +Z axis aligns with its forward viewing direction. Consequently, the upward direction corresponds to the -Y axis.
@@
As discussed in the main paper’s conclusion, a current limitation of our BevSplat method, which renders Bird’s-Eye View (BEV) perspectives based on 3D Gaussian Splatting [kerbl20233d], is its computational speed compared to Inverse Perspective Mapping (IPM). For instance, on an NVIDIA RTX 4090 GPU, BevSplat requires 14 ms to generate a single BEV image. In contrast, IPM, which utilizes direct linear interpolation, can achieve this in 4ms. This performance disparity affects the overall inference speed of our model. Therefore, a significant direction for our future work is the exploration of faster and more compact Gaussian representations to address this bottleneck and enhance real-time applicability.